# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 5.8 Kinetic Viscous Heating
As discussed above for the case of magnetic pumping, in weakly collisional plasmas, low-frequency turbulent fluctuations that generate variations in the magnetic field magnitude will lead to a double-adiabatic (CGL) (Chew et al., 1956) evolution of the plasma that preserves to lowest order the particle magnetic moment $\mu_{m}$ and the action integral of the parallel bounce motion $J_{\parallel}$ (Lichko et al., 2017; Montag, 2018; Arzamasskiy et al., 2023). This dynamics leads to the development of anisotropies in the velocity distribution of the plasma particles exhibiting temperature and pressure anisotropies with $p_{\parallel}\neq p_{\perp}$. These deviations from local thermodynamic equilibrium can trigger rapidly growing kinetic instabilities that can disrupt the linear wave dynamics (Squire et al., 2016, 2017b, 2017a) and mediate nonlocal transfer of turbulent energy directly from the large driving scales to kinetic length scales (Kunz et al., 2018; Arzamasskiy et al., 2023), as illustrated by the red arrow in Fig. 1(b). The effect of these kinetic instabilities on the turbulent dynamics has recently been modeled as an effective viscosity (Arzamasskiy et al., 2023).
@@
To properly describe the kinetic physics of turbulence in temperature anisotropic plasmas, it is necessary to employ the full set of plasma and turbulence parameters in Table 2. The relevant kinetic instabilities driven by anisotropies in the ion velocity distribution are the four ion temperature anisotropy instabilities: (i) the parallel (or whistler) firehose instability (Kennel, 1966; Gary et al., 1976) and (ii) the Alfvén (or oblique) firehose instability (Hellinger & Matsumoto, 2000), which both are relevant to plasmas with an ion temperature anisotropy sufficiently less than unity $A_{i}=T_{\perp,i}/T_{\parallel,i}<1$ and with a parallel ion plasma beta $\beta_{i}>1$; and (iii) the mirror instability (Vedenov & Sagdeev, 1958; Tajiri, 1967; Southwood & Kivelson, 1993) and (iv) the ion cyclotron instability (Gary et al., 1976), which both are relevant to plasmas with an ion temperature anisotropy sufficiently greater than unity $A_{i}=T_{\perp,i}/T_{\parallel,i}>1$ for any value of the parallel ion plasma beta $\beta_{i}$. When the plasma exceeds a threshold value of the proton temperature anisotropy (Boris:1977; Matteini:2006; Hellinger:2008; Klein:2015), these instabilities can tap the free energy associated with the anisotropic proton temperature, driving electromagnetic fluctuations and ultimately reducing the temperature anisotropy, thereby moving the plasma back toward a state of marginal stability.
@@
Recent theoretical and numerical studies have highlighted the fact that, at sufficiently high parallel ion plasma beta $\beta_{i}\gg 1$, the marginal stability boundary for these kinetic ion temperature anisotropy instabilities asymptotes towards isotropy, so the double-adiabatic evolution of turbulent fluctuations can easily exceed these boundaries and trigger these rapidly growing instabilities. Large-scale Alfvén waves in the MHD limit at $k_{\perp}\rho_{i}\ll 1$ yield only perpendicular magnetic field perturbations $\delta B_{\perp}\neq 0$ with $\delta B_{\parallel}=0$, so the variation in the magnetic field magnitude $\delta B$ is negligible in the limit of small amplitude waves $\delta B_{\perp}/B_{0}\ll 1$. However, at sufficiently large amplitudes $\delta B_{\perp}/B_{0}\rightarrow 1$, these fluctuations can lead to significant variation in the magnetic field magnitude, as made clear by (39). Thus, at sufficiently high values of parallel ion plasma beta $\beta_{i}\gg 1$, the physics of linearly polarized [^11] Note that circularly polarized Alfvénic fluctuations, which have no magnetic field magnitude variation, do not suffer this kinetic-instability-driven interruption (Squire et al., 2016). Alfvén waves can be “interrupted” by these kinetic ion temperature anisotropy instabilities (Squire et al., 2016, 2017b, 2017a).
@@
Recent theoretical and numerical work has extended this concept to consider the effect of kinetic instabilities on the turbulent cascade under conditions of high parallel ion plasma beta $\beta_{i}\gg 1$, finding that these instabilities can nonlocally transfer energy directly to kinetic length scales (Arzamasskiy et al., 2023), as illustrated by the red arrow in Fig. 1(b). This new channel of energy transfer can compete with the traditional local transfer of energy to smaller scales of the turbulent cascade that is mediated by nonlinear interactions, as illustrated by the black arrows in Fig. 1(a). The effect of these kinetic instabilities on the large-scale turbulent dynamics has been quantified as an effective viscosity by Arzamasskiy et al. (2023). Here we denote this kinetic-instability-mediated viscosity as the kinetic viscosity [^12] This kinetic viscosity terminology for kinetic-instability-mediated viscosity should not be confused with the kinematic viscosity defined in hydrodynamic fluids. to distinguish it from the usual definition of viscosity that arises through Coulomb collisions among the plasma particles.
@@
These ion kinetic instabilities act to limit the ion pressure anisotropy in the plasma to values corresponding to marginal stability, given by $\Delta p_{i}/p_{i}\lesssim 1/\beta_{i}$, where $\Delta p_{i}\equiv p_{\perp i}-p_{\parallel i}$ and $p_{i}\equiv(2p_{\perp i}+p_{\parallel i})/3$ (Kunz et al., 2014; Melville et al., 2016; Arzamasskiy et al., 2023). Using this marginal stability scaling with $\beta_{i}$, one can estimate an effective collision frequency $\nu_{\mbox{eff}}\sim\beta_{i}\omega(\delta B_{\perp}/B_{0})^{2}$ (Arzamasskiy et al., 2023), where $\omega$ is the linear wave frequency at the scale where the dynamics generate that largest pressure anisotropy, and thus have the maximum growth rate. Since the turbulent fluctuation amplitudes generally decrease monotonically with scale, the kinetic temperature anisotropy instabilities will be driven most strongly by turbulent fluctuations at the driving scale. As with the case for magnetic pumping by compressible fluctuations that is dominated by the driving scale dyanmics, here we must use the three driving parameters $(k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0})$ in Table 2, rather than $k_{0}\rho_{i}$, to estimate the turbulent damping rate by the kinetic viscosity.
@@
Using the effective collision frequency $\nu_{\mbox{eff}}$ above, we can estimate the ratio of the damping rate due to kinetic viscous heating to the wave frequency at the driving scales, obtaining
@@
$\frac{\gamma_{VH}}{\omega}\sim\beta_{i}\left(\chi_{0}\frac{k_{\parallel 0}}{k_{\perp 0}}\right)^{2},$ (42)
@@
In summary, the kinetic viscous heating due to ion temperature anisotropy instabilities arising from the nonthermal velocity distributions driven by the large-scale turbulent motions occurs in weakly collisional turbulent plasmas with parallel ion plasma beta $\beta_{i}\gg 1$. The kinetic viscous damping rate depends on the parallel ion plasma beta $\beta_{i}$, the nonlinearity parameter $\chi_{0}$, and the anisotropy of the turbulent driving $k_{\perp 0}/k_{\parallel 0}$. Note that the scaling used to derive the form of the effective collision frequency $\nu_{\mbox{eff}}$ depends on the scaling of the marginal stability boundaries that are generally calculated in quiescient plasmas in the absence of turbulence. The scaling of marginal stability boundaries in the presence of strong plasma turbulence may differ, although evidence from spacecraft observations (Hellinger et al., 2006; Bale et al., 2009; Maruca et al., 2011) and from hybrid kinetic ion and fluid electron simulations of temperature anisotropic plasma turbulence (Kunz et al., 2014; Bott et al., 2021; Arzamasskiy et al., 2023) appear to suggest that any modifications of these boundaries due to the presence of turbulence are relatively small.
@@ 5.9 Collisionless Magnetic Reconnection
Under the B06 theory for the scaling of MHD turbulence, the phenomenon of dynamic alignment (Boldyrev, 2006) is predicted to lead to the development of current sheets with a width $w\sim 1/k_{i}$ to thickness $a\sim 1/k_{\perp}$ ratio in the plane perpendicular to the magnetic field given by $w/a=k_{\perp}/k_{i}=[(k_{\perp}\rho_{i})/(k_{0}\rho_{i})]^{1/4}$, according to the scaling in (2) for $\alpha=1$. The geometry of the current sheets expected to develop in plasma turbulence is illustrated in Fig. 8, where for sufficiently small perpendicular scales relative to the driving scale, $k_{\perp}\rho_{i}\gg k_{0}\rho_{i}$, the current sheets have a parallel extent along the mean magnetic field $\mathbf{B}_{0}$, given by $l\sim 1/k_{\parallel}$, and a scale separation of thickness to width to length obeying $a\ll w\ll l$; the corresponding three components of the wavenumber have the complementary scaling $k_{\parallel}\ll k_{i}\ll k_{\perp}$. In Fig. 8, the wavelength of a mode that is unstable to the collisionless tearing instability is illustrated by $\lambda\sim 1/k$ (red). Note that the minimum value of this tearing unstable wavenumber $k$ is equal to the intermediate wavenumber $k_{i}$ corresponding to the width of the current sheet, $w\sim 1/k_{i}$. The maximum value of $k$ assumed in calculations of the linear collisionless tearing instability (Loureiro & Boldyrev, 2017a; Mallet et al., 2017b) must be much smaller than the perpendicular wavenumber $k_{\perp}$ characterizing the current sheet thickness ($a\sim 1/k_{\perp}$), thereby yielding $k/k_{\perp}=ka\ll 1$. Thus, the range of $k$ is given by
@@
Figure 8: Diagram of the current sheet geometry generated self-consistently in plasma turbulence in the MHD regime at $k_{\perp}\rho_{i}\lesssim 1$, with the scalngs of $(k_{\perp},k_{i},k_{\parallel})$ given by the B06 scaling. The current sheet has thickness $a\sim k_{\perp}$ and width $w\sim 1/k_{i}$ in the perpendicular plane and length $l\sim 1/k_{\parallel}$ along the equilibirium magnetic field $\mathbf{B}_{0}$. The wavelength of the mode unstable to the collisionless tearing instability is given by $\lambda\sim 1/k$ (red), and must fall within the bounds determined by the current sheet scaling, yielding an ordering $k_{i}\lesssim k\ll k_{\perp}$.
Note that, in turbulence obeying the B06 scaling, the self-consistent generation of current sheets arises only at sufficiently small perpendicular scales, $k_{\perp}\rho_{i}\gg k_{0}\rho_{i}$. Therefore, the ratio of the perpendicular magnetic field perturbation to the mean magnetic field will be small, $\delta B_{\perp}/B_{0}=(k_{0}\rho_{i}/k_{\perp}\rho_{i})^{1/4}\ll 1$, so the reconnection that arises in plasma turbulence necessarily falls in the limit of strong-guide-field magnetic reconnection, as illustrated in Fig. 8 (green). An implication of this fact is that, if numerical simulations are employed to assess the importance of magnetic reconnection in turbulence, it is critical that those simulations be modeled in three spatial dimensions (3D). Previous studies of 2D vs. 3D plasma turbulence have shown that 3D simulations exhibit a significantly faster turbulent cascade rate (Li et al., 2016), so enabling full 3D evolution is likely to be essential to determine the competition between the turbulent cascade and magnetic reconnection.
The tearing instability growth rate typically increases with an increasing perpendicular aspect ratio $w/a$ ratio of the current sheets. If this ratio reaches a sufficiently large value, it was recently recognized that the growth rate of the tearing instability could exceed the turbulent cascade rate at that scale, $\gamma/\omega_{nl}\gtrsim 1$, triggering a reconnection flow that would disrupt the turbulent dynamics, deviating from the scaling relations summarized in §3 and potentially affecting the damping of turbulent cascade (Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017b, a; Mallet et al., 2017b, a; Walker et al., 2018).
@@
Here we adopt the B06 scaling for the development of current sheets to determine the dependence of magnetic reconnection as a turbulent damping mechanism on the fundamental plasma and turbulence parameters for the isotropic temperature case in Table 3. This calculation follows previous investigations that assess the conditions under which collisionless magnetic reconnection may disrupt the turbulent cascade and thereby impact how the turbulent energy is channeled into particle energy (Loureiro & Boldyrev, 2017a; Mallet et al., 2017b). Our goal here is to express the results from these previous studies in the dimensionless parameters for the isotropic temperature case proposed in Table 3.
The growth rate of the collisionless tearing instabilty, which is the instability that initiates the process of magnetic reconnection under the weakly collisional conditions typical of space and astrophysical plasmas, has been derived in the limit of very low electron plasma beta $\beta_{e}\sim\mu^{-1}$ (Zocco & Schekochihin, 2011), where we convert $\beta_{e}=\beta_{i}/\tau$. The physics of the collisionless tearing instabilty depends on three length scales: (i) the thickness of the current sheet $a\sim 1/k_{\perp}$, assumed here to occur on MHD scales with $k_{\perp}\rho_{i}\ll 1$; (ii) the ion sound Larmor radius $\rho_{s}\equiv\rho_{i}/(2\tau)^{1/2}$ where two fluid effects begin to lead to differences between the ion and electron dynamics; and (iii) the inner boundary layer scale $\delta_{in}$ where the magnetic flux can be unfrozen from the electron fluid flow. We will restrict our analysis to $\beta_{e}\gg\mu^{-1}$, such that the flux unfreezing arises from electron inertia (Zocco & Schekochihin, 2011; Loureiro & Boldyrev, 2017a; Mallet et al., 2017b), occuring approximately at the electron inertial length scale $d_{e}\equiv c/\omega_{pe}=\rho_{i}/(\beta_{i}\mu)^{1/2}$. These scales are ordered by $\delta_{in}\sim d_{e}\ll\rho_{s}\ll a$. Together, these restrictions limit us to moderately small plasma beta,
@@
The tearing instabilty depends on the tearing mode instability parameter $\Delta^{\prime}$, determined by the outer-region solution (at MHD scales $k_{\perp}\rho_{i}\ll 1$) (Zocco & Schekochihin, 2011). This parameter can be expressed over two limiting regimes in a dimensionless form (Loureiro & Boldyrev, 2017a) as
@@
This equation shows that $\Delta^{\prime}\delta_{in}$ increases as the perpendicular wavenumber $k_{\perp}\rho_{i}$ increases, corresponding to a thinner current sheet, or as the unstable wavenumber $k\rho_{i}$ decreases, requiring to a wider current sheet. Since the width $w\sim 1/k_{i}$ of the current sheets in B06 scaling theory governed by (2) bounds the unstable wavenumber $k$ from below by the intermediate wavenumber, $k>k_{i}$, the maximum value of instability parameter $\Delta^{\prime}\delta_{in}$ increases with the perpendicular aspect ratio of the current sheets $w/a=k_{\perp}/k_{i}\gg 1$. Note also that $\Delta^{\prime}\delta_{in}\propto\beta_{i}^{-1/3}$, so that lower values of ion plasma beta lead to a larger instability parameter.

### v2

@@ 5.8 Kinetic Viscous Heating
As discussed above for the case of magnetic pumping, low-frequency turbulent fluctuations that generate variations in the magnetic field magnitude will lead to a double-adiabatic (CGL) (Chew et al., 1956) evolution of weakly collisional plasmas that preserves to lowest order the particle magnetic moment $\mu_{m}$ and the action integral of the parallel bounce motion $J_{\parallel}$ (Lichko et al., 2017; Montag, 2018; Arzamasskiy et al., 2023). This dynamics leads to the development of anisotropies in the velocity distribution of the plasma particles, yielding temperature and pressure anisotropies with $p_{\parallel}\neq p_{\perp}$. These deviations from local thermodynamic equilibrium can trigger rapidly growing kinetic instabilities that can disrupt the linear wave dynamics (Squire et al., 2016, 2017a, 2017b) and mediate the nonlocal transfer of turbulent energy directly from the large driving scales to kinetic length scales (Kunz et al., 2018; Arzamasskiy et al., 2023), as illustrated by the red arrow in Fig. 1(b). The effect of these kinetic instabilities on the turbulent dynamics has recently been modeled as an effective “viscosity” (Arzamasskiy et al., 2023).
Here we use the term “kinetic viscous heating” to describe the transfer of energy by this effective viscosity mediated by temperature anisotropy instabilities (Arzamasskiy et al., 2023) to avoid confusion with the usual viscous heating arising from microscopic Coulomb collisions. We note that, although this effect can be grossly modeled as an effective viscosity that leads to diffusion of the bulk fluid velocity field, its properties are entirely distinct from the standard viscosity in fluid dynamics mediated by microscopic collisions between neutral particles (Chapman & Cowling, 1970) or by microscopic Coulomb collisions between charged particles in a plasma under strongly collisional conditions (Spitzer, 1962; Grad, 1963; Braginskii, 1965). Standard viscosity is a linear mechanism that may be modeled mathematically by a Laplacian differential operator in the fluid momentum equation, and thus its effect increases monotonically with $k^{2}$ in Fourier space; this kinetic viscosity is a nonlinear mechanism operating only under weakly collisional conditions that strongly depends on the value of the ion plasma beta $\beta_{i}$, and its effect diminishes at higher wavenumber $k$ with the monotonically decreasing amplitudes of the turbulent fluctuations. Determining how this kinetic viscosity may tap some fraction of the energy from the turbulent fluctuations at the large, driving scales and nonlocally transfer that energy directly to fluctuations at the small, ion kinetic length scales remains an open line of investigation.
@@
To properly describe the kinetic physics of turbulence in temperature anisotropic plasmas, it is necessary to employ the full set of plasma and turbulence parameters in Table 3. The relevant kinetic instabilities driven by anisotropies in the ion velocity distribution are the four ion temperature anisotropy instabilities: (i) the parallel (or whistler) firehose instability (Kennel, 1966; Gary et al., 1976) and (ii) the Alfvén (or oblique) firehose instability (Hellinger & Matsumoto, 2000), which are both relevant to plasmas with an ion temperature anisotropy sufficiently less than unity $A_{i}=T_{\perp,i}/T_{\parallel,i}<1$ and with a parallel ion plasma beta $\beta_{\parallel,i}>1$; and (iii) the mirror instability (Vedenov & Sagdeev, 1958; Tajiri, 1967; Southwood & Kivelson, 1993) and (iv) the ion cyclotron instability (Gary et al., 1976), which are both relevant to plasmas with an ion temperature anisotropy sufficiently greater than unity $A_{i}=T_{\perp,i}/T_{\parallel,i}>1$ for any value of the parallel ion plasma beta $\beta_{\parallel,i}$. When the plasma exceeds a threshold value of the ion temperature anisotropy (Boris & Manheimer, 1977; Matteini et al., 2006; Hellinger & Trávníček, 2008; Klein & Howes, 2015), these instabilities can tap the free energy associated with the anisotropic ion temperature, driving electromagnetic fluctuations and ultimately reducing the temperature anisotropy, thereby moving the plasma back toward a state of marginal stability.
Hellinger et al. (2006) has compiled values for the marginal stability boundaries for these four ion temperature anisotropy instabilities over the $(\beta_{\parallel,i},T_{\perp,i}/T_{\parallel,i})$ plane. Using a linear Vlasov-Maxwell dispersion relation solver with bi-Maxwellian equilibrium ion velocity distributions in a fully ionized, hydrogenic (proton and electron) plasma, the marginal stability boundary is determined by calculating the complex eigenfrequency $\omega(\mathbf{k})$ for a fixed $\beta_{\parallel,i}$ over all possible wavevectors $\mathbf{k}$. The ion temperature anisotropy $T_{\perp,i}/T_{\parallel,i}$ is then varied until the most unstable wavevector has a growth rate of $|\gamma|/\Omega_{i}=10^{-3}$, thus establishing the instability criterion. For the four ion temperature anisotropy instabilities, the corresponding instability criteria on the $(\beta_{\parallel,i},T_{\perp,i}/T_{\parallel,i})$ plane are generally well fit by an expression of the form
$\frac{T_{\perp,i}}{T_{\parallel,i}}=1+\frac{a}{(\beta_{\parallel,i}-\beta_{0})^{b}}$ (44)
where $a,\ b,$ and $\beta_{0}$ are unique values for each of the four instabilities, computed by Hellinger et al. (2006) and found here in Table 6. These instability thresholds are plotted in Fig. 8, and observational studies of measured intervals in the solar wind generally show that the measurements are more or less constrained by these threshold limits on the $(\beta_{\parallel,i},T_{\perp,i}/T_{\parallel,i})$ plane (Hellinger et al., 2006; Bale et al., 2009; Maruca et al., 2011).
Figure 8: Marginal linear stability thresholds for the four ion temperature anisotropy instabilities using (44) and the parameters in Table 6.
Instability $a$ $b$ $\beta_{0}$
Ion Cyclotron 0.43 0.42 -0.0004
Parallel Firehose -0.47 0.53 0.59
Alfvén Firehose -1.4 1.0 -0.11
Mirror 0.77 0.76 -0.016
Table 6: Instability Threshold Parameters for a maximum growth rate $|\gamma|/\Omega_{i}=10^{-3}$ from Hellinger et al. (2006).
@@
Recent theoretical and numerical studies have highlighted the fact that, at sufficiently high parallel ion plasma beta $\beta_{\parallel,i}\gg 1$, the marginal stability boundary for these kinetic ion temperature anisotropy instabilities asymptotes towards isotropy, as shown in Fig. 8, so the double-adiabatic evolution of turbulent fluctuations can easily exceed these boundaries and trigger these rapidly growing instabilities. Large-scale Alfvén waves in the MHD limit at $k_{\perp}\rho_{i}\ll 1$ yield only perpendicular magnetic field perturbations $\delta B_{\perp}\neq 0$ with $\delta B_{\parallel}=0$, so the variation in the magnetic field magnitude $\delta B$ is negligible in the limit of small amplitude waves $\delta B_{\perp}/B_{0}\ll 1$. However, at sufficiently large amplitudes $\delta B_{\perp}/B_{0}\rightarrow 1$, these fluctuations can lead to significant variation in the magnetic field magnitude, as made clear by (41). Thus, at sufficiently high values of parallel ion plasma beta $\beta_{\parallel,i}\gg 1$, the physics of linearly polarized [^14] Note that circularly polarized Alfvénic fluctuations, which have no magnetic field magnitude variation, do not suffer this kinetic-instability-driven interruption (Squire et al., 2016). Alfvén waves can be “interrupted” by these kinetic ion temperature anisotropy instabilities (Squire et al., 2016, 2017a, 2017b).
@@
Further examination of the effect of kinetic instabilities on the turbulent cascade under conditions of high parallel ion plasma beta, $\beta_{\parallel,i}\gg 1$, has found that these instabilities can nonlocally transfer energy directly to kinetic length scales (Arzamasskiy et al., 2023), as illustrated by the red arrow in Fig. 1(b). This new channel of energy transfer can compete with the traditional local transfer of energy to smaller scales of the turbulent cascade that is mediated by nonlinear interactions, as illustrated by the black arrows in Fig. 1(a). The effect of these kinetic instabilities on the large-scale turbulent dynamics has been quantified as an effective viscosity by Arzamasskiy et al. (2023). Here we denote this kinetic-instability-mediated viscosity as the kinetic viscosity [^15] This kinetic viscosity terminology for kinetic-instability-mediated viscosity should not be confused with the kinematic viscosity defined in hydrodynamic fluids. to distinguish it from the usual definition of viscosity that arises through Coulomb collisions among the charged plasma particles.
@@
These ion kinetic instabilities act to limit the ion pressure anisotropy in the plasma to values corresponding to marginal stability, given by $\Delta p_{i}/p_{i}\lesssim 1/\beta_{\parallel,i}$, where $\Delta p_{i}\equiv p_{\perp i}-p_{\parallel i}$ and $p_{i}\equiv(2p_{\perp i}+p_{\parallel i})/3$ (Kunz et al., 2014; Melville et al., 2016; Arzamasskiy et al., 2023). Using this marginal stability scaling with $\beta_{\parallel,i}$, one can estimate an effective collision frequency (Arzamasskiy et al., 2023)
$\nu_{\mbox{eff}}\sim\beta_{\parallel,i}\omega(\delta B_{\perp}/B_{0})^{2},$ (45)
where $\omega$ is the linear wave frequency at the scale where the dynamics generate the largest pressure anisotropy, thus leading to the maximum instability growth rate. Since the turbulent fluctuation amplitudes generally decrease monotonically with scale, the kinetic temperature anisotropy instabilities will be driven most strongly by turbulent fluctuations at the driving scale. As with the case for magnetic pumping by compressible fluctuations that is dominated by the driving scale dynamics, here we must use the three driving parameters $(k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0})$ in Table 3, rather than $k_{0}\rho_{i}$, to estimate the turbulent damping rate by the kinetic viscosity.
@@
Using the effective collision frequency $\nu_{\mbox{eff}}$ above as an estimate of the damping rate by this kinetic viscous heating $\gamma_{VH}$, we can calculate the ratio of the damping rate arising from kinetic viscous heating to the wave frequency at the driving scale, obtaining
@@
$\frac{\gamma_{VH}}{\omega}\sim\beta_{\parallel,i}\left(\chi_{0}\frac{k_{\parallel 0}}{k_{\perp 0}}\right)^{2},$ (46)
@@
In summary, the kinetic viscous heating due to ion temperature anisotropy instabilities arising from the nonthermal velocity distributions driven by the large-scale turbulent motions occurs in weakly collisional turbulent plasmas with parallel ion plasma beta $\beta_{\parallel,i}\gg 1$. The kinetic viscous damping rate depends on the parallel ion plasma beta $\beta_{\parallel,i}$, the nonlinearity parameter $\chi_{0}$, and the anisotropy of the turbulent driving $k_{\perp 0}/k_{\parallel 0}$. It is important to keep in mind that the scaling used to derive the form of the effective collision frequency $\nu_{\mbox{eff}}$ depends on the scaling of the marginal stability boundaries that are generally calculated in quiescent plasmas in the absence of turbulence. The scaling of the marginal stability boundaries—such as those quantified by (44) and the coefficients in Table 6—may differ in the presence of strong plasma turbulence, although evidence from spacecraft observations (Hellinger et al., 2006; Bale et al., 2009; Maruca et al., 2011) and from hybrid kinetic ion and fluid electron simulations of temperature anisotropic plasma turbulence (Kunz et al., 2014; Bott et al., 2021; Arzamasskiy et al., 2023) appear to suggest that any modifications of these boundaries due to the presence of turbulence are relatively small.
@@ 5.9 Collisionless Magnetic Reconnection
Under the B06 theory for the scaling of MHD turbulence, the phenomenon of dynamic alignment (Boldyrev, 2006) is predicted to lead to the development of current sheets with a width $w\sim 1/k_{i}$ to thickness $a\sim 1/k_{\perp}$ ratio in the plane perpendicular to the magnetic field given by $w/a=k_{\perp}/k_{i}=[(k_{\perp}\rho_{i})/(k_{0}\rho_{i})]^{1/4}$, according to the scaling in (2) for $\alpha=1$. The geometry of the current sheets expected to develop in plasma turbulence is illustrated in Fig. 9, where for sufficiently small perpendicular scales relative to the driving scale, $k_{\perp}\rho_{i}\gg k_{0}\rho_{i}$, the current sheets have a parallel extent along the mean magnetic field $\mathbf{B}_{0}$, given by $l\sim 1/k_{\parallel}$, and a scale separation of thickness to width to length obeying $a\ll w\ll l$; the corresponding three components of the wave vector have the complementary scaling $k_{\parallel}\ll k_{i}\ll k_{\perp}$. In Fig. 9, the wavelength of a mode that is unstable to the collisionless tearing instability is illustrated by $\lambda\sim 1/k$ (red). Note that the minimum value of this tearing unstable wavenumber $k$ is equal to the intermediate wavenumber $k_{i}$ corresponding to the width of the current sheet, $w\sim 1/k_{i}$. The maximum value of $k$ assumed in calculations of the linear collisionless tearing instability (Loureiro & Boldyrev, 2017a; Mallet et al., 2017b) must be much smaller than the perpendicular wavenumber $k_{\perp}$ characterizing the current sheet thickness ($a\sim 1/k_{\perp}$), thereby yielding $k/k_{\perp}=ka\ll 1$. Thus, the range of the unstable wavenumber $k$ is given by
@@
Figure 9: Diagram of the current sheet geometry generated self-consistently in plasma turbulence in the MHD regime at $k_{\perp}\rho_{i}\lesssim 1$, with the scalings of $(k_{\perp},k_{i},k_{\parallel})$ given by the B06 scaling. The current sheet has thickness $a\sim 1/k_{\perp}$ and width $w\sim 1/k_{i}$ in the perpendicular plane and length $l\sim 1/k_{\parallel}$ along the equilibrium magnetic field $\mathbf{B}_{0}$. The wavelength of the mode unstable to the collisionless tearing instability is given by $\lambda\sim 1/k$ (red), and must fall within the bounds determined by the current sheet scaling, yielding an ordering $k_{i}\lesssim k\ll k_{\perp}$.
Note that, in turbulence obeying the B06 scaling, the self-consistent generation of current sheets arises only at sufficiently small perpendicular scales relative to the driving scale, $k_{\perp}\rho_{i}\gg k_{0}\rho_{i}$. Therefore, the ratio of the perpendicular magnetic field perturbation to the mean magnetic field will be small, $\delta B_{\perp}/B_{0}=(k_{0}\rho_{i}/k_{\perp}\rho_{i})^{1/4}\ll 1$, so the reconnection that arises in plasma turbulence necessarily falls in the limit of strong-guide-field magnetic reconnection, as illustrated in Fig. 9 (green). An implication of this fact is that, if numerical simulations are employed to assess the importance of magnetic reconnection in turbulence, it is critical that those simulations be modeled in three spatial dimensions (3D). Previous studies of 2D vs. 3D plasma turbulence have shown that 3D simulations exhibit a significantly faster turbulent cascade rate (Li et al., 2016), so enabling full 3D evolution is likely to be essential to determine the competition between the turbulent cascade and magnetic reconnection.
The tearing instability growth rate typically increases with an increasing perpendicular aspect ratio $w/a$ ratio of the current sheets. If this ratio reaches a sufficiently large value, it was recently recognized that the growth rate of the tearing instability could exceed the turbulent cascade rate at that scale, $\gamma/\omega_{nl}\gtrsim 1$, triggering a reconnection flow that would disrupt the turbulent dynamics, deviating from the scaling relations summarized in §3. Such changes in the scaling of the turbulent fluctuations may potentially lead to differences in which physical mechanisms dominate the damping of the turbulent cascade at small scales (Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017b, a; Mallet et al., 2017b, a; Walker et al., 2018).
@@
Here we adopt the B06 scaling for the development of current sheets to determine the dependence of magnetic reconnection as a turbulent damping mechanism on the fundamental plasma and turbulence parameters for the isotropic temperature case in Table 4. This calculation follows previous investigations that assess the conditions under which collisionless magnetic reconnection may disrupt the turbulent cascade and thereby impact how the turbulent energy is channeled into particle energy (Loureiro & Boldyrev, 2017a; Mallet et al., 2017b). Our goal here is to express the results from these previous studies in terms of the dimensionless parameters for the isotropic temperature case proposed in Table 4.
The growth rate of the collisionless tearing instability, which is the instability that initiates the process of magnetic reconnection under the weakly collisional conditions typical of space and astrophysical plasmas, has been derived in the limit of very low electron plasma beta $\beta_{e}\sim\mu^{-1}$ (Zocco & Schekochihin, 2011), where we convert $\beta_{e}=\beta_{i}/\tau$. The physics of the collisionless tearing instability depends on three length scales: (i) the thickness of the current sheet $a\sim 1/k_{\perp}$, assumed here to occur on MHD scales with $k_{\perp}\rho_{i}\ll 1$; (ii) the ion sound Larmor radius $\rho_{s}\equiv\rho_{i}/(2\tau)^{1/2}$ where two fluid effects begin to lead to differences between the ion and electron dynamics; and (iii) the inner boundary layer scale $\delta_{in}$ where the magnetic flux can be unfrozen from the electron fluid flow. We will restrict our analysis to $\beta_{e}\gg\mu^{-1}$, such that the flux unfreezing arises from electron inertia (Zocco & Schekochihin, 2011; Loureiro & Boldyrev, 2017a; Mallet et al., 2017b), occurring approximately at the electron inertial length scale $d_{e}\equiv c/\omega_{pe}=\rho_{i}/(\beta_{i}\mu)^{1/2}$. These scales are ordered by $\delta_{in}\sim d_{e}\ll\rho_{s}\ll a$. Together, these restrictions limit us to moderately small plasma beta,
@@
The tearing instability depends on the tearing mode instability parameter $\Delta^{\prime}$, determined by the outer-region solution (at MHD scales $k_{\perp}\rho_{i}\ll 1$) (Zocco & Schekochihin, 2011). This parameter can be expressed over two limiting regimes in a dimensionless form (Loureiro & Boldyrev, 2017a) as
@@
This equation shows that $\Delta^{\prime}\delta_{in}$ increases as the perpendicular wavenumber $k_{\perp}\rho_{i}$ increases, corresponding to a thinner current sheet, or as the unstable wavenumber $k\rho_{i}$ decreases, requiring to a wider current sheet. Since the width $w\sim 1/k_{i}$ of the current sheets in B06 scaling theory governed by (2) bounds the unstable wavenumber $k$ from below by the intermediate wavenumber, $k>k_{i}$, the maximum value of instability parameter $\Delta^{\prime}\delta_{in}$ increases with the perpendicular aspect ratio of the current sheets $w/a=k_{\perp}/k_{i}\gg 1$. Note also that $\Delta^{\prime}\delta_{in}\propto\beta_{i}^{-1/3}$, so that lower values of ion plasma beta also lead to a larger instability parameter.